# Vision-Language Models for Medical Report Generation and Visual Question Answering: A Review

### v1

@@ 4 Medical Vision-Language Datasets
  • MIMIC-CXR-JPG [Joh+19] is a pre-processed variant of the MIMIC-CXR [Joh+19a] dataset. In this version, the original $377,110$ images are converted into compressed JPG format. The $227,827$ reports associated with these images are enriched with labels for various common pathologies. The labels are derived from the analysis of the impression, findings, or final sections of the radiology reports, facilitated by the use of NegBio [Pen+17] and CheXpert (Chest eXpert) [Irv+19] tools.
@@
  • MIMIC-NLE [Kay+22] dataset is specifically designed for the task of generating natural language explanations (NLEs) to justify predictions made on medical images, particularly in the context of thoracic pathologies and chest X-ray findings. The dataset consists of $38,003$ image-NLE pairs or $44,935$ image-diagnosis-NLE triplets, acknowledging instances where a single NLE may explain multiple diagnoses. NLEs are extracted from MIMIC-CXR [Joh+19a] radiology reports. The dataset exclusively considers X-ray views from front-to-back (anteroposterior, AP) and back-to-front (posteroanterior, PA). All NLEs come with diagnosis and evidence (for a diagnosis) labels. The dataset is split into the training set with $37,016$ images, a test set with $273$ images, and a validation set with $714$ images.
@@
  • CXR-PRO (CXR with Prior References Omitted) [RCR22] dataset is derived from MIMIC-CXR [Joh+19a]. The dataset consists of $374,139$ free-text radiology reports containing only the impression sections. It also incorporates associated chest radiographs; however, the radiology reports and chest X-rays are not paired. This dataset is designed to mitigate the problem of hallucinated references to prior reports often generated by radiology report generation ML models. The omission of prior references in this dataset aims to provide a cleaner and more reliable dataset for radiology RG.
@@
  • IU-Xray ((Indiana University chest X-rays) dataset, also known as the Open-I [Dem+15] dataset, is accessible through the National Library of Medicine’s Open-i service. The dataset originates from two hospital systems within the Indiana Network for Patient Care database. This dataset comprises $7,470$ DICOM chest X-rays paired with $3,955$ associated radiology reports. The reports typically include sections such as indications, findings, and impressions, and they are manually annotated using MeSH and RadLex (Radiology Lexicon) codes to represent clinical findings and diagnoses. Throughout this review, we will refer to the dataset interchangeably as IU-Xray and Open-I, maintaining consistency with the nomenclature used in related literature.
@@
  • MedICaT (Medical Images, Captions, and Textual Reference) [Sub+20] dataset contains $217,060$ figures from $131,410$ open-access PMC papers focused on radiology images and other medical imagery types. Excluding figures from ROCO [Pel+18], the dataset integrates inline references from the S2ORC (Semantic Scholar Open Research Corpus) [Lo+20] corpus, establishing connections between references and corresponding figures. Additionally, the inline references to ROCO figures are provided separately. MedICaT also contains $7,507$ subcaption-subfigure pairs with annotations derived from $2,069$ compound figures.
@@
  • PMC-OA (PubMedCentral’s OpenAccess) [Lin+23] dataset comprises $1.65$ M image-caption pairs, derived from PMC papers. It encompasses a variety of diagnostic procedures, including common ones such as ultrasound, MRI, PET, radioisotope, and rarer procedures like mitotic and fMRI. Additionally, the dataset covers a broad spectrum of diseases, with induced cataracts, ear diseases, and low vision being among the most frequently represented conditions.
@@
  • MS-CXR [Boe+22] dataset contains image bounding box labels paired with radiology findings, annotated and verified by two board-certified radiologists. The dataset consists of $1,162$ image-text pairs of bounding boxes and corresponding text descriptions. The annotations cover 8 different cardiopulmonary radiological findings and are extracted from MIMIC-CXR [Joh+19a] and REFLACX (Reports and Eye-tracking data For Localization of Abnormalities in Chest X-rays) [Big+22] (based on MIMIC-CXR) datasets. The findings include atelectasis, cardiomegaly, consolidation, edema, lung opacity, pleural effusion, pneumonia, and pneumothorax.
@@
  • SLAKE (Semantically-Labeled Knowledge-Enhanced) is a English-Chinese bilingual dataset [Liu+21]. It contains $642$ images, including $12$ diseases and $39$ organs of the whole body. Each image is meticulously annotated with two types of visual information: masks for semantic segmentation and bounding boxes for object detection. The dataset includes a total of $14,028$ QA pairs, categorized into vision-only or knowledge-based types and labeled accordingly, encompassing both open- and closed-ended questions. Moreover, SLAKE incorporates $5,232$ medical knowledge triplets in the form of $<head,relation,tail>$, where $head$ and $tail$ denote entities (e.g., organ, disease), and $relation$ signify the relationship between these entities (e.g., function, treatment). An illustrative example of such a triplet is $<pneumonia,location,lung>$.
@@
  • VQA-RAD [Lau+18] dataset contains $104$ head axial single-slice CTs or MRIs, $107$ chest x-rays, and $104$ abdominal axial CTs. The images are meticulously chosen from MedPix, an open-access online medical image database, ensuring each image corresponds to a unique patient. Furthermore, every selected image has an associated caption and is deliberately devoid of any radiology markings. Every caption provides details about the imaging plane, modality, and findings generated and reviewed by expert radiologists. Also, VQA-RAD contains $3,515$ QA pairs, with an average of 10 questions per image. Among them, $,51$ 515 are free-form questions and answers, allowing for unrestricted inquiry. Additionally, $733$ pairs involve rephrased questions and answers, introducing linguistic diversity. Another $1,267$ pairs are framed, featuring questions presented in a structured format, offering consistency and systematic evaluation. Additionally, QA pairs are split into $637$ open-ended and $878$ closed-ended types. Within the closed-ended group, a predominant focus is on yes/no questions.
@@
  • PathVQA [He+20] is a dataset that encompasses $4,998$ pathology images accompanied by a total of $32,799$ QA pairs derived from these images. The images are sourced from pathology books: “Textbook of Pathology” and “Basic Pathology”, and the digital library “Pathology Education Informational Resource”. Out of all QA pairs, $16,465$ are of the open-ended type, while the remaining pairs are of the closed-ended yes/no type. On average, each image is associated with $6.6$ questions, which cover a broad spectrum of visual contents, encompassing aspects such as color, location, appearance, shape, etc.
@@
  • VQA-Med 2019 [Aba+19] dataset contains $4,200$ radiology images obtained from MedPix, an open-access online medical image database, and $15,292$ QA pairs. The training set consists of $3,200$ images and $12,792$ QA pairs, with each image having 3 to 4 associated questions. The validation set includes $500$ images and $2,000$ QA pairs, and the test set comprises $500$ images and $500$ QA pairs. The questions are mainly about modality, imaging plane, organ system, and abnormality.
@@
  • VQA-Med 2020 [Aba+20] dataset contains $5,000$ radiology images obtained from MedPix, an open-access online medical image database, and $5,000$ QA pairs. The training set consists of $4,000$ images and $4,000$ QA pairs. The validation set comprises $500$ images and $500$ QA pairs, and the test set includes $500$ images and $500$ QA pairs. The questions are focused on abnormalities present in the images. Additionally, the dataset contains radiology images and questions for the Visual Question Generation (VQG) task. The training set consists of $780$ images and $2,156$ associated questions. The validation set comprises $141$ images with $164$ questions, and the test set includes $80$ images.
@@
  • VQA-Med 2021 [Ion+21] dataset contains $5,500$ radiology images obtained from MedPix, an open-access online medical image database, and $5,500$ QA pairs. The training set consists of $4,500$ images and $4,5000$ QA pairs. The validation set comprises $500$ images and $500$ QA pairs, and the test set includes $500$ images and $500$ QA pairs. The questions are focused on abnormalities present in the images. Similarly to VQA-Med 2019, the dataset also contains radiology images and questions for the VQG task. The validation set comprises $85$ images with $200$ questions, and the test set includes $100$ images.
@@
  • EndoVis 2017 (Endoscopic Vision) [All+19] dataset contains 5 robotic surgery videos (two videos with 8 frames each, one with 18, one with 14, and one with 39 frames) from the MICCAI (Medical Image Computing and Computer Assisted Interventions) Endoscopic Vision 2017 Challenge. It also includes 472 QA pairs with bounding box annotations. These QA pairs are carefully crafted to involve specific inquiries related to the surgical procedure. Examples of questions include queries such as ”What is the state of prograsp forceps?” and “Where is the large needle driver located?”. The inclusion of bounding box annotations enhances the dataset’s utility for tasks such as object detection or answer localization.
@@
  • EndoVis 2018 [All+20] dataset contains 14 robotic surgery videos ($2,007$ frames in total) from the MICCAI Endoscopic Vision 2018 Challenge. It also includes $11,783$ QA pairs regarding organs, surgical tools, and organ-tool interactions. When the question is about organ-tool interactions, the bounding box will contain both the organ and the tool.
5 Vision-Language Model Evaluation
@@
5.1 Evaluation Metrics for Report Generation
The prevalent benchmark datasets for medical RG are MIMIC-CXR [Joh+19a] and Open-I [Dem+15]. For more information on these datasets, see Section 4. Several metrics are used to evaluate the effectiveness of VLMs on RG tasks. The more frequently used metrics are outlined below.
  • BLEU (Bilingual Evaluation Understudy) [Pap+02] score was originally designed for machine translation evaluation, but it has been adapted for RG and even VQA in a modified form. BLEU provides a quantitative measure of how well the machine-generated text aligns with human-generated reference text. First, the precision of different n-grams, which are consecutive sequences of n words, is calculated using the formula:
@@ 5.1 Evaluation Metrics for Report Generation
    where overlapping n-grams refer to n-grams in the model-generated text that share common elements with at least one n-gram in the reference text. To ensure the precision score remains robust and is not disproportionately affected by repeated n-grams in the model-generated text, a modification known as clipping is often introduced. This process involves capping the count of each n-gram in the model-generated text to a maximum count. This maximum count is determined by the highest count observed in any single reference text for the same n-gram. The final BLUE-n score is defined as:
@@
    where $BP$ is a brevity penalty calculated as
    $BP=\begin{cases}1&\text{if }c\geq r\\ e^{(1-r/c)}&\text{if }c<r\end{cases}$
    where $c$ is the length of the model-generated text, and $r$ is the length of the reference text. It is common to use $n=4$. The BLEU score ranges from 0 to 1, where a higher score suggests better agreement with the reference text. The overall BLUE score of the model is the average of BLUE scores for each pair of reports.
  • ROUGE (Recall-Oriented Understudy for Gisting Evaluation) [Lin04] is a set of metrics that evaluate the overlap between the model-generated text and human-generated reference text. ROUGE-n assesses the overlap of n-grams between model-generated text and reference text, and it is defined as:
@@
    ROUGE-L focuses on measuring the longest common subsequence between model-generated text $Y$ and reference text $X$ and it is calculated using the formula:
@@
    where $R=\frac{LCS(X,Y)}{m}$, $P=\frac{LCS(X,Y)}{n}$, $m$ is the length of $X$, $n$ is the length of $Y$, $LCS(X,Y)$ is the length of a longest common subsequence of $X$ and $Y$, and $\beta$ is a parameter that depends on the specific task and the relative importance of precision (P) and recall (R). There are other ROUGE score variants. The ROUGE scores range from 0 to 1, where higher scores indicate similarity between the model-generated text and the reference text. For each ROUGE variant, the overall score of the model is the average of scores for each instance.
  • METEOR (Metric for Evaluation of Translation with Explicit ORrdering) [BL05] is an evaluation metric designed to be more forgiving than some other metrics and takes into account the fluency and meaning of the generated text. The METEOR score is computed as follows:
    $METEOR=\frac{10\times P\times R}{R+9\times P}(1-Penalty),$
@@
    $R=\frac{\#\text{\emph{overlapping 1-grams}}}{\#\text{\emph{1-grams in a reference text}}},\ P=\frac{\#\text{\emph{overlapping 1-grams}}}{\#\text{\emph{1-grams in a model-generated text}}},$
    $Penalty=\frac{1}{2}\times\left(\frac{\#chunks}{\#\text{\emph{overlapping 1-grams}}}\right)^{3},$
    $chunks$ are groups of adjacent 1-grams in the model-generated text that overlap with adjacent 1-grams in the reference text. The METEOR score ranges from 0 to 1, with higher scores indicating better alignment between the model-generated text and the reference text. The overall METEOR score of a model is the average of scores for each instance.
  • Perplexity [Hao+20] measures the average uncertainty of a model in predicting each word in a text. The formula for perplexity is defined as:
    $Perplexity=e^{-\frac{1}{n}\sum_{k=1}^{n}\ln P(w_{k}|w_{1},w_{2},\dots,w_{k-1})},$
    where $n$ is the total number of words in the text. The value of the perplexity metric can range from 1 to positive infinity, and lower values signify a more accurate and confident model in capturing the language patterns within the given text.
  • BERTScore [Zha+20a] metric was initially designed for evaluating models that use BERT [Dev+19] embeddings. However, it can also leverage other word embeddings to evaluate the similarity between model-generated and reference text. The BERTScore of a single text pair is calculated according to the formula:
    $BERTScore=\frac{2\times P\times R}{P+R},$
@@
  • RadGraph F1 [Yu+23] is a novel metric that measures overlap in clinical entities and relations extracted from radiology reports. The RadGraph F1 score is computed in the following way. First, the RadGraph model maps model-generated and reference reports into graph representations with clinical entities represented as nodes and their relations as edges between them. Second, the number of nodes that match between the two graphs based on clinical entity text and labels (entity type) is determined. Third, the number of edges that match between the two graphs based on their start and end entities and labels (relation type) is calculated. Lastly, the F1 score is separately computed for clinical entities and relations, and then the RadGraph F1 score for a report pair is the average of these two scores. The overall model performance is determined by averaging RadGraph F1 scores across all report pairs.
  • Human evaluation plays a crucial role in assessing VLMS’s quality in medical RG. Human evaluation can be performed for RG in various ways. For instance, in [Jeo+23] expert radiologists evaluate the performance of the X-REM model in the RG task as follows. Initially, each report is segmented into lines, and radiologists assign scores to each line based on five error categories. These scores reflect the severity of errors, with higher values indicating more severe errors. Two metrics are utilized to obtain a comprehensive measure of the overall severity of errors in a report. Maximum Error Severity (MES) represents the highest score across all lines in the report, while Average Error Severity (AES) is calculated by averaging the scores across all lines in the report. According to radiologists, $18\%$ of model-generated reports received an MES score of 00, while $24\%$ received an AES score of 00.
The next few metrics are designed for classification evaluation, and RG can be viewed as such a task. In [Moo+22], [Lee+23], and [Pel+23], these metrics are computed based on the 14 labels obtained from applying the CheXpert [Irv+19] or CheXbert [Smi+20] labeler to the reference reports as well as the model-generated reports. In this context, reports bearing accurate diagnosis labels are categorized as positive, while those with inaccurate labels are regarded as negative. The following metrics are also called clinical efficacy metrics.
@@
  • Precision is a metric that measures the accuracy of positive predictions made by a model. The precision score is calculated by considering the ratio of true positive predictions to the total number of instances that the model predicted as positive:
    $Presicion=\frac{\text{\emph{True Positivies}}}{\text{\emph{True Positivies + False Positivies}}}.$
@@
    $Recall=\frac{\text{\emph{True Positives}}}{\text{\emph{True Positives + False Negatives}}}.$
@@
  • F1 metric assesses the overall model’s performance by balancing precision and recall into a single value. The F1 score is defined as:
    $F1=\frac{2\times\text{\emph{Precision}}\times\text{\emph{Recall}}}{\text{\emph{Precision + False Recall}}}.$
@@
5.2 Evaluation Metrics for Visual Question Answering
The common benchmark datasets for medical VQA include VQA-RAD [Lau+18], SLAKE [Liu+21], and PathVQA [He+20]. For more information on these datasets, see Section 4. Various metrics are available for VQA evaluation, and many of those used for RG can also be applied to VQA. To avoid redundancy with already mentioned metrics, only a few are highlighted below.
  • Accuracy is a fundamental metric for gauging overall model correctness in VQA evaluation. It is determined by calculating the proportion of correctly predicted answers to the total number of questions. Sometimes, the average accuracy is computed by applying the model to various testing datasets, providing a comprehensive assessment of its performance across diverse scenarios.

### v2

@@ 4.1.3 MIMIC-CXR-JPG
MIMIC-CXR-JPG [johnson2019mimiccxrjpg] is a pre-processed variant of the MIMIC-CXR dataset [johnson2019mimiccxr]. In this version, the original $377,110$ images are converted into compressed JPG format. The $227,827$ reports associated with these images are enriched with labels for various common pathologies. The labels are derived from the analysis of the impression, findings, or final sections of the radiology reports, facilitated by the use of NegBio [peng2017negbio] and CheXpert (Chest eXpert) [irvin2019chexpert] tools.
4.1.4 MIMIC-NLE
@@
MIMIC-NLE dataset is specifically designed for the task of generating natural language explanations (NLEs) to justify predictions made on medical images, particularly in the context of thoracic pathologies and chest X-ray findings [kayser2022mimicnle]. The dataset consists of $38,003$ image-NLE pairs or $44,935$ image-diagnosis-NLE triplets, acknowledging instances where a single NLE may explain multiple diagnoses. NLEs are extracted from MIMIC-CXR [johnson2019mimiccxr] radiology reports. The dataset exclusively considers X-ray views from front-to-back (anteroposterior, AP) and back-to-front (posteroanterior, PA). All NLEs come with diagnosis and evidence (for a diagnosis) labels. The dataset is split into the training set with $37,016$ images, a test set with $273$ images, and a validation set with $714$ images.
4.1.5 CXR with Prior References Omitted (CXR-PRO)
@@
CXR-PRO dataset is derived from MIMIC-CXR [johnson2019mimiccxr]. The dataset consists of $374,139$ free-text radiology reports containing only the impression sections [ramesh2022cxrredone]. It also incorporates associated chest radiographs; however, the radiology reports and chest X-rays are not paired. This dataset is designed to mitigate the problem of hallucinated references to prior reports often generated by radiology report generation ML models. The omission of prior references in this dataset aims to provide a cleaner and more reliable dataset for radiology RG.
4.1.6 Indiana University chest X-rays (IU-Xray)
@@
IU-Xray dataset, also known as the Open-I dataset, is accessible through the National Library of Medicine’s Open-i service [demnerfushman2015openi]. The dataset originates from two hospital systems within the Indiana Network for Patient Care database. This dataset comprises $7,470$ DICOM chest X-rays paired with $3,955$ associated radiology reports. The reports typically include sections such as indications, findings, and impressions, and they are manually annotated using MeSH and RadLex (Radiology Lexicon) codes to represent clinical findings and diagnoses. Throughout this review, we will refer to the dataset interchangeably as IU-Xray and Open-I, maintaining consistency with the nomenclature used in related literature.
4.1.7 Medical Images, Captions, and Textual References (MedICaT)
@@
MedICaT dataset contains $217,060$ figures from $131,410$ open-access PMC papers focused on radiology images and other medical imagery types [subramanian2020medicat]. Excluding figures from ROCO [pelka2018roco], the dataset integrates inline references from the S2ORC (Semantic Scholar Open Research Corpus) [lo2020s2orc] corpus, establishing connections between references and corresponding figures. Additionally, the inline references to ROCO figures are provided separately. MedICaT also contains $7,507$ subcaption-subfigure pairs with annotations derived from $2,069$ compound figures.
4.1.8 PubMedCentral’s OpenAccess (PMC-OA)
@@
PMC-OA dataset comprises $1.65$ M image-caption pairs, derived from PMC papers [lin2023pmcclip]. It encompasses a variety of diagnostic procedures, including common ones such as ultrasound, MRI, PET, and radioisotope, and rarer procedures like mitotic and fMRI. Additionally, the dataset covers a broad spectrum of diseases, with induced cataracts, ear diseases, and low vision being among the most frequently represented conditions.
4.1.9 MS-CXR
@@
MS-CXR dataset contains image bounding box labels paired with radiology findings, annotated and verified by two board-certified radiologists [boecking2022mscxr]. The dataset consists of $1,162$ image-text pairs of bounding boxes and corresponding text descriptions. The annotations cover 8 different cardiopulmonary radiological findings and are extracted from MIMIC-CXR [johnson2019mimiccxr] and REFLACX (Reports and Eye-tracking data For Localization of Abnormalities in Chest X-rays) [lanfredi2022reflacx] (based on MIMIC-CXR) datasets. The findings include atelectasis, cardiomegaly, consolidation, edema, lung opacity, pleural effusion, pneumonia, and pneumothorax.
4.1.10 Semantically-Labeled Knowledge-Enhanced (SLAKE)
@@
SLAKE is an English-Chinese bilingual dataset [liu2021slake]. It contains $642$ images, including $12$ diseases and $39$ organs of the whole body. Each image is meticulously annotated with two types of visual information: masks for semantic segmentation and bounding boxes for object detection. The dataset includes a total of $14,028$ QA pairs, categorized into vision-only or knowledge-based types and labeled accordingly, encompassing both open- and closed-ended questions. Moreover, SLAKE incorporates $5,232$ medical knowledge triplets in the form of $<head,relation,tail>$, where $head$ and $tail$ denote entities (e.g., organ, disease), and $relation$ signify the relationship between these entities (e.g., function, treatment). An illustrative example of such a triplet is $<$pneumonia, location, lung$>$.
4.1.11 VQA-RAD
@@
VQA-RAD dataset contains $104$ head axial single-slice CTs or MRIs, $107$ chest x-rays, and $104$ abdominal axial CTs [lau2018vqarad]. The images are meticulously chosen from MedPix, an open-access online medical image database, ensuring each image corresponds to a unique patient. Furthermore, every selected image has an associated caption and is deliberately devoid of any radiology markings. Every caption provides details about the imaging plane, modality, and findings generated and reviewed by expert radiologists. Also, VQA-RAD contains $3,515$ QA pairs, with an average of 10 questions per image. Among them, $,51$ 515 are free-form questions and answers, allowing for unrestricted inquiry. Additionally, $733$ pairs involve rephrased questions and answers, introducing linguistic diversity. Another $1,267$ pairs are framed, featuring questions presented in a structured format, offering consistency and systematic evaluation. Additionally, QA pairs are split into $637$ open-ended and $878$ closed-ended types. Within the closed-ended group, a predominant focus is on yes/no questions.
4.1.12 PathVQA
@@
PathVQA is a dataset that encompasses $4,998$ pathology images accompanied by a total of $32,799$ QA pairs derived from these images [he2020pathvqa]. The images are sourced from pathology books: “Textbook of Pathology” and “Basic Pathology”, and the digital library “Pathology Education Informational Resource”. Out of all QA pairs, $16,465$ are of the open-ended type, while the remaining pairs are of the closed-ended yes/no type. On average, each image is associated with $6.6$ questions, which cover a broad spectrum of visual contents, encompassing aspects such as color, location, appearance, shape, etc.
4.1.13 VQA-Med 2019
@@
VQA-Med 2019 dataset contains $4,200$ radiology images obtained from MedPix, an open-access online medical image database, and $15,292$ QA pairs [abacha2019vqamed19]. The training set consists of $3,200$ images and $12,792$ QA pairs, with each image having 3 to 4 associated questions. The validation set includes $500$ images and $2,000$ QA pairs, and the test set comprises $500$ images and $500$ QA pairs. The questions are mainly about modality, imaging plane, organ system, and abnormality.
4.1.14 VQA-Med 2020
@@
VQA-Med 2020 dataset contains $5,000$ radiology images obtained from MedPix, an open-access online medical image database, and $5,000$ QA pairs [abacha2020vqamed20]. The training set consists of $4,000$ images and $4,000$ QA pairs. The validation set comprises $500$ images and $500$ QA pairs, and the test set includes $500$ images and $500$ QA pairs. The questions are focused on abnormalities present in the images. Additionally, the dataset contains radiology images and questions for the Visual Question Generation (VQG) task. The training set consists of $780$ images and $2,156$ associated questions. The validation set comprises $141$ images with $164$ questions, and the test set includes $80$ images.
4.1.15 VQA-Med 2021
@@
VQA-Med 2021 dataset contains $5,500$ radiology images obtained from MedPix, an open-access online medical image database, and $5,500$ QA pairs [ionescu2021vqamed21]. The training set consists of $4,500$ images and $4,5000$ QA pairs. The validation set comprises $500$ images and $500$ QA pairs, and the test set includes $500$ images and $500$ QA pairs. The questions are focused on abnormalities present in the images. Similarly to VQA-Med 2019, the dataset also contains radiology images and questions for the VQG task. The validation set comprises $85$ images with $200$ questions, and the test set includes $100$ images.
4.1.16 Endoscopic Vision (EndoVis) 2017
@@
EndoVis 2017 dataset contains 5 robotic surgery videos (two videos with 8 frames each, one with 18, one with 14, and one with 39 frames) from the MICCAI (Medical Image Computing and Computer Assisted Interventions) Endoscopic Vision 2017 Challenge [allan2019endovis17]. It also includes 472 QA pairs with bounding box annotations. These QA pairs are carefully crafted to involve specific inquiries related to the surgical procedure. Examples of questions include queries such as ”What is the state of prograsp forceps?” and “Where is the large needle driver located?”. The inclusion of bounding box annotations enhances the dataset’s utility for tasks such as object detection or answer localization.
4.1.17 EndoVis 2018
@@
EndoVis 2018 dataset contains 14 robotic surgery videos ($2,007$ frames in total) from the MICCAI Endoscopic Vision 2018 Challenge [allan2020endovis18]. It also includes $11,783$ QA pairs regarding organs, surgical tools, and organ-tool interactions. When the question is about organ-tool interactions, the bounding box will contain both the organ and the tool.
4.2 VLM Evaluation Metrics
@@
4.2.1 Evaluation Metrics for Report Generation
The prevalent benchmark datasets for medical RG are MIMIC-CXR [johnson2019mimiccxr] and Open-I [demnerfushman2015openi]. For more information on these datasets, see Section 4.1. Several metrics are used to evaluate the effectiveness of VLMs on RG tasks. The more frequently used metrics are outlined below.
Bilingual Evaluation Understudy (BLEU)
The BLEU score was originally designed for machine translation evaluation, but it has been adapted for RG and even VQA in a modified form. BLEU provides a quantitative measure of how well the machine-generated text aligns with human-generated reference text [papineni2002bleu]. First, the precision of different n-grams, which are consecutive sequences of $n$ words, is calculated using the formula:
@@ 4.2.1 Evaluation Metrics for Report Generation
where ‘overlapping n-grams’ refer to n-grams in the model-generated text that share common elements with at least one n-gram in the reference text. To ensure the precision score remains robust and is not disproportionately affected by repeated n-grams in the model-generated text, a modification known as clipping is often introduced. This process involves capping the count of each n-gram in the model-generated text to a maximum count. This maximum count is determined by the highest count observed in any single reference text for the same n-gram. The final BLEU-n score is defined as:
@@
In eq. 2, $BP$ is referred to as the brevity penalty and is calculated as:
$BP=\begin{cases}1&\text{if }c\geq r\\ e^{(1-r/c)}&\text{if }c<r,\end{cases}$ (3)
where $c$ is the length of the model-generated text, and $r$ is the length of the reference text. It is common to use $n=4$. The BLEU score ranges from 0 to 1, where a higher score suggests better agreement with the reference text. The overall BLEU score of the model is the average of BLEU scores for each pair of reports.
Recall-Oriented Understudy for Gisting Evaluation (ROUGE)
ROUGE is a set of metrics that evaluate the overlap between the model-generated text and human-generated reference text [lin2004rouge]. ROUGE-n assesses the overlap of n-grams between model-generated text and reference text, and it is defined as:
@@
ROUGE-L focuses on measuring the longest common subsequence between model-generated text $Y$ and reference text $X$, and it is calculated using the following relationship:
@@
where $R={LCS(X,Y)}/{m}$, $P={LCS(X,Y)}/{n}$, $m$ is the length of $X$, $n$ is the length of $Y$, $LCS(X,Y)$ is the length of a longest common subsequence of $X$ and $Y$, and $\beta$ is a parameter that depends on the specific task and the relative importance of precision (P) and recall (R). There are other ROUGE score variants. The ROUGE scores range from 0 to 1, where higher scores indicate similarity between the model-generated text and the reference text. For each ROUGE variant, the overall score of the model is the average of scores for each instance.
Metric for Evaluation of Translation with Explicit ORrdering (METEOR)
METEOR is an evaluation metric designed to be more forgiving than some other metrics and takes into account the fluency and meaning of the generated text [banarjee2005meteor]. The METEOR score is computed as follows:
$\text{METEOR}=\frac{10\times P\times R}{R+9\times P}(1-\text{Penalty})$ (6)
@@
$R$ =#$=\frac{\#\text{overlapping 1-grams}}{\#\text{1-grams in a reference text}},$ (7)
$P$ =#$=\frac{\#\text{overlapping 1-grams}}{\#\text{1-grams in a model-generated text}},$ (8)
Penalty $=\frac{1}{2}\times\left(\frac{\#\text{chunks}}{\#\text{overlapping 1-grams}}\right)^{3},$ (9)
and $chunks$ are groups of adjacent 1-grams in the model-generated text that overlap with adjacent 1-grams in the reference text. The METEOR score ranges from 0 to 1, with higher scores indicating better alignment between the model-generated text and the reference text. The overall METEOR score of a model is the average of scores for each instance.
Perplexity
Perplexity measures the average uncertainty of a model in predicting each word in a text [hao2020preplexity]. The formula for perplexity is defined as:
$\text{Perplexity}=\exp\left({-\frac{1}{n}\sum_{k=1}^{n}\ln P(w_{k}|w_{1},w_{2},\dots,w_{k-1})}\right),$ (10)
where $n$ is the total number of words in the text. The value of the perplexity metric can range from 1 to +$\infty$+ ∞, and lower values signify a more accurate and confident model in capturing the language patterns within the given text.
BERTScore
BERTScore metric was initially designed for evaluating models that use BERT [devlin2019bert] embeddings [zhang2020bertscore]. However, it can also leverage other word embeddings to evaluate the similarity between model-generated and reference text. The BERTScore of a single text pair is calculated according to the relationship:
$\text{BERTScore}=\frac{2\times P\times R}{P+R},$ (11)
@@
RadGraph F1
RadGraph F1 is a novel metric that measures overlap in clinical entities and relations extracted from radiology reports [yu2023radgraph]. The RadGraph F1 score is computed in the following way. First, the RadGraph model maps model-generated and reference reports into graph representations with clinical entities represented as nodes and their relations as edges between them. Second, the number of nodes that match between the two graphs based on clinical entity text and labels (entity type) is determined. Third, the number of edges that match between the two graphs based on their start and end entities and labels (relation type) is calculated. Lastly, the F1 score is separately computed for clinical entities and relations, and then the RadGraph F1 score for a report pair is the average of these two scores. The overall model performance is determined by averaging RadGraph F1 scores across all report pairs.
Human evaluation
Human evaluation plays a crucial role in assessing VLMS’s quality in medical RG. Human evaluation can be performed for RG in various ways. For instance, in [jeong2023xrem], expert radiologists evaluate the performance of the X-REM model in the RG task as follows. Initially, each report is segmented into lines, and radiologists assign scores to each line based on five error categories. These scores reflect the severity of errors, with higher values indicating more severe errors. Two metrics are utilized to obtain a comprehensive measure of the overall severity of errors in a report. Maximum Error Severity (MES) represents the highest score across all lines in the report. In contrast, Average Error Severity (AES) is calculated by averaging the scores across all lines in the report. According to radiologists, $18\%$ of model-generated reports received an MES score of 00, while $24\%$ received an AES score of 00.
Additional Evaluation Metrics for Report Generation
The next few metrics are designed for classification evaluation, and RG can be viewed as such a task. In [moon2022medvill], [lee2023unixgen], and [pellegrini2023radialog], these metrics are computed based on the 14 labels obtained from applying the CheXpert [irvin2019chexpert] or CheXbert [smit2020chexbert] labeler to the reference reports as well as the model-generated reports. In this context, reports bearing accurate diagnosis labels are categorized as positive, while those with inaccurate labels are regarded as negative. The following metrics are also called clinical efficacy metrics.
@@
  • Precision measures the accuracy of positive predictions made by a model. The precision score is calculated by considering the ratio of true positive predictions to the total number of instances that the model predicted as positive:
    Presicion=True PositiviesTrue Positivies + False $\text{Presicion}=\frac{\text{True Positivies}}{\text{True Positivies + False Positivies}}.$ (12)
@@
    Recall=True PositivesTrue Positives + False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives +False Negatives}}.$ (13)
@@
  • F1 Score assesses the overall model’s performance by balancing precision and recall into a single value. The F1 score is defined as:
[…]
@@
4.2.2 Evaluation Metrics for VQA
The common benchmark datasets for medical VQA include VQA-RAD [lau2018vqarad], SLAKE [liu2021slake], and PathVQA [he2020pathvqa]. For more information on these datasets, see Section 4.1. Various metrics are available for VQA evaluation, and many of those used for RG can also be applied to VQA. To avoid redundancy with already mentioned metrics, only a few are highlighted below.
Accuracy
Accuracy is a fundamental metric for gauging overall model correctness in VQA evaluation. It is determined by calculating the proportion of correctly predicted answers to the total number of questions. Sometimes, the average accuracy is computed by applying the model to various testing datasets, providing a comprehensive assessment of its performance across diverse scenarios. For a detailed comparison of accuracies among different medical VLMs discussed in Section 4.3, refer to Table 3.
Exact Match